Daytime

Nighttime

DARK LGSW: 20150226_021930

We set the FLAO cube rotator to an angle of 208° (best position to minimize vignetting on LUCI according to July 2014 alignment).

We take a set of LUCI images at different cube rotator angles to verify this position (LUCI2.20150206.0135.fits is the one corresponding to 208°).

20:00: we point NGC2419, preset active sent from LUCI.

Nominal board position to center the stages on the LUCI rotator axis are: -74mm, -50mm.

PBL with pointing the new software doesn't work, we revert to the previous build.

Zenith operation

The LGS are far off their nominal position. We point the telescope to zenith and we go on the telescope to check their position.

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes
20150226_044159 0, 0, 0 NA NA F F Light off

LM1 A1 A2 A3
Abs 15024 19888 22356
Rel -2179 -546 -4418
We acquire LM1 to QC interaction matrix since it doesn't need any NGS.

LM1-QC REC: 20150226_054500

NGC2419

We point back to NGC2419: the pointing model is not working anymore, we have to offset by hand the LM1 by +20theta and -20phi.

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes
20150226_050141 0, 0, 0 NA NA F F
We try to optimize the LGS spot moving the astigmatism plate and taking snapshots for analysis.

22:30: actual positions

Focus -150k
Corr 800
Rot 6500
1:50: we had clouds up to this time. Now seeing is about 0.7".

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes
20150226_104301 0.3, -0, 0 NA NA F F QC only, ~800 counts
20150226_104549 1, -0, 0 NA NA T F
20150226_104932 1,-0.2, 0 NA NA F F QC + LGSW
20150226_105513 1,-0.4,-0.01 NA NA T T
20150226_105622 1,-0.4,-0.1 NA NA T T
Clouds are thinning, ~1400 counts on QC, m=13.5 on PYR BIN4 with 50-50 on FW1.

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes Plot
20150226_113706 2,-0.3, -0 133.448 20150224_093700 T F 20150226_113706-20150226_121019_modalplot.png
20150226_120815 1, -0, -0 333.656 20150224_093700 F F
20150226_120843 1, -0, -0 344.357 20150224_093700 F F
20150226_120903 1, -0, -0 389.661 20150224_093700 F F
20150226_120942 1, -0, -0 409.276 20150224_093700 F F
20150226_121019 1, -0, -0 401.744 20150224_093700 F F
Now we run the LGSW at 500Hz because we lack light:

TN AO Gain Res. WF [nm rms] AO rec JCL PCL Notes Plot
20150226_121612 -0.4,-0.3,-0.1 1082.29 20141202_125700 F F 20150226_121612-20150226_122516_modalplot.png
20150226_121725 -0.4,-0.5,-0.2 1235.13 20141202_125700 F F
20150226_122043 -0.4,-0.6,-0.4 1116.63 20141202_125700 F F
20150226_122148 -0.4,-0.6,-0.4 1041.38 20141202_125700 F F
20150226_122232 -0.4,-0.7,-0.5 1060.23 20141202_125700 F F
20150226_122318 -0.4,-0.7,-0.5 993.38 20141202_125700 F F
20150226_122516 -1,-0.6,-0.4 1259.43 20141202_125700 F F
20150226_122722 0, 0, 0 NA NA F F -
20150226_124052 0, 0, 0 404.084 20150224_093700 F F 20150226_124052-20150226_124442_modalplot.png
20150226_124358 1, -0, -0 273.213 20150224_093700 F F
20150226_124442 1, -0, -0 297.642 20150224_093700 F F
20150226_124951 1,-0.1,-0.01 NA NA F F -

20150226_104301-20150226_125218_LgsFluxVsTime.png 20150226_104301-20150226_125218_LgsFwhmVsTime.png

PockelsCellsTransmission test

Test performed at the end of the night because we could not keep closed the LGS loop on more than 36 modes... Sweep between 0.1 and 2.0us in 7 steps with Internal sources.

TN Sup Plot
''20150226_130745, '20150226_131053' B=68, Y=122, R=109 20150226_130745_pcTransmission.jpg
-- MarcoBonaglia - 05 Mar 2015

Topic revision: r3 - 17 Mar 2015, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback